The sum of Two numbers is 39.Twice the first number and plus 3 times the second number is 101.What are the numbers?

Answer by ursuchacrass(5)

The sum of Two numbers is 39.Twice the first number and plus 3 times the second
number is 101.What are the numbers?
Solution
let the two numbers be x and y
x+y=39.......equation 1
2x+3y=101....equation 2

get the value of y from the equation one....
y=-x+39 then substitute y to equation two...
2x+3(-x+39)=101
2x-3x+117=101....group like terms
2x-3x=101-117
-x=-16
x=16
substitute the value of x to first equation
x+y=39
16+y=39
y=39-16
y=23


Checking....
2x+3y=101
2(16)+3(23)=101
32+69=101
101=101

it's easy...
(madali lang ito)
From matthew..huhuhu